Determine the critical​ t-scores for each of the conditions below.

​a)

​one-tail test,

alphaαequals=0.01,

and

nequals=38

​b)

​one-tail test,

alphaαequals=0.005,

and

nequals=10

​c)

​two-tail test,

alphaαequals=0.05

and

nequals=25

​d)

​two-tail test,

alphaα=0.10

and

n=13

Click here to view page 1 of the Student's t-distribution table.

LOADING...

Click here to view page 2 of the Student's t-distribution table.

LOADING...

​a) The critical​ t-score(s) for a​ one-tail test, where

alphaα=0.01​,

and

n=38

​is(are)

nothing.

​(Round to three decimal places as needed. Use a comma to separate answers as​ needed.)

​b) The critical​ t-score(s) for a​ one-tail test, where

alphaα=0.005

and

n=10

​is(are)

nothing.

​(Round to three decimal places as needed. Use a comma to separate answers as​ needed.)

​c) The critical​ t-score(s) for a​ two-tail test, where

alphaα=0.05​,

and

n=2

​is(are)

nothing.

​(Round to three decimal places as needed. Use a comma to separate answers as​ needed.)

​d) The critical​ t-score(s) for a​ two-tail test, where

alphaαequals=0.10

and

n=13

​is(are)

nothing.

​(Round to three decimal places as needed. Use a comma to separate answ
Answer #1

Solution :

Given that,

(a)

sample size = n = 38

Degrees of freedom = df = n - 1 = 38 - 1 = 37

= 0.01

t ,df = t 0.01,37 = 2.431

critical t score = 2.431

(b)

sample size = n = 10

Degrees of freedom = df = n - 1 = 10 - 1 = 9

= 0.005

t ,df = t 0.005,9 = 3.250

critical t score = 3.250

(c)

sample size = n = 2

Degrees of freedom = df = n - 1 = 2 - 1 = 1

= 0.05

/ 2 = 0.05 / 2 = 0.025

t/2 ,df = t 0.025,1 = -/+12.706

critical t score = -12.706 , +12.706

(d)

sample size = n = 13

Degrees of freedom = df = n - 1 = 13 - 1 = 12

= 0.10

/ 2 = 0.10 / 2 = 0.05

t/2 ,df = t 0.05,12 = -/+1.782

critical t score = -1.782 , +1.782